OpenStudy (anonymous):

Represent the number 2.5 as a percent and as a reduced fraction.

OpenStudy (anonymous):

I assume 2.5 is the number. as a 2.5*100% = 250% as a fraction it is easy to see that this is 250/100 (b/c is = 2.5) dividing 250 and 100 by 50, you get 5/2 (= 2.5)